OpenStudy (anonymous):

find the slopes of the curve y^4=y^2-x^2 at (sqrt(3)/4,-1/2) and (-sqrt(3)/4,1/2)

OpenStudy (amistre64):

take the derivative and toss in your x and y values to determine slope as usual

OpenStudy (amistre64):

D( y^4=y^2-x^2 ) 4y^3 y' =2y y' -2x x' 2y^3 y' =y y' -x x' and since x' is just dx/dx = 1 2y^3 y' =y y' -x algebra the y' all by itself and fill in the given point values for x and y